OpenStudy (anonymous):

Evaluate LIM 3x^2 - (7x+1) sqrt root (x) + 5 . ------------------------- X---> 1 (X - 1)

OpenStudy (aravindg):

do you know l hospital rule?

OpenStudy (anonymous):

Were not in there yet just. the basics Rationalizing/Factorization

OpenStudy (aravindg):

i really couldnt understand ur numerator..can u please use equation editor to make it readable?

OpenStudy (anonymous):

Oh wait sorry

OpenStudy (anonymous):

LIM X->1

OpenStudy (anonymous):

3x^2 - (7x+1) square root of (X) +5 ----------------------------------- X-1

OpenStudy (aravindg):

\[\sqrt{x+5}??\]

OpenStudy (callisto):

Is this the question?\[\lim_{x \to 1}\frac{3x^2-(7x+1)\sqrt{x}+5}{x-1}\]

OpenStudy (anonymous):

Yeah

OpenStudy (callisto):

A bit complicated, forgive me. \[\lim_{x \to 1}\frac{3x^2-(7x+1)\sqrt{x}+5}{x-1}\]\[=\lim_{x \to 1}\frac{3x^2-(7x+1)\sqrt{x}+5}{(\sqrt{x}-1)(\sqrt{x}+1)}\]\[=\lim_{x \to 1}\frac{3x^2-7x^\frac{3}{2}-\sqrt{x}+5}{(\sqrt{x}-1)(\sqrt{x}+1)}\]To make the substitution works, we need to get \(\sqrt{x}-1\) cancelled. So, we have to factorize the numerator with \(\sqrt{x}-1\) as a factor. Hmm.. Can you do it?

OpenStudy (anonymous):

I''ll try But how did it end up in conjugating ? how did it became square root of X in the denominator

OpenStudy (callisto):

\[x-1 = (\sqrt{x})^2 -1 =(\sqrt{x}-1)(\sqrt{x}+1)\]

OpenStudy (callisto):

I think in this case, we don't need conjugate :|

OpenStudy (callisto):

We only need factorization.